# Non-coherent evolution of closed weakly interacting system leads to equidistribution of probabilities of microstates

### v5

@@ III Non-coherent evolution
If $P_{l_{N}}$ is a smooth function of energy at the point of interest, in a short interval we can approximate it as a constant, so summation over $l_{N}$ gives the number of states inside the energy region – $N$ (Fig. 4b):
$\frac{\pi}{\hbar\Delta E}\sum_{l_{T},l_{N}}|V_{kl}|^{2}P_{l_{T}}=\frac{\pi N}{\hbar\Delta E}\sum_{l_{T}}|V_{kl}|^{2}P_{l_{T}}=\frac{2\pi}{\hbar}\sum_{l_{T}}|V_{kl}|^{2}P_{l_{T}}\nu(E)$ (47)
where $\nu(E)=N/(2\Delta E)$ is the dencity of states. Multiplier 2 arises because $\Delta E$ is the absolute value of the deviation of energy from $E_{l}$ and the energy in the region can be less than $E_{l}$ and higher than $E_{l}$ so the length of the region is $2\Delta E$.
We can reason differently. We perform the same subsitution $\Delta t=\pi\hbar/\Delta E$ as in (46), and approximate the sum as an integral:
$\frac{1}{\hbar^{2}}\sum_{l}|V_{kl}|^{2}\Delta tP_{l}=\frac{2\pi}{\hbar}C\int dl\frac{\overline{\chi}_{\Delta E}}{\Delta E}|V_{kl}|^{2}P_{l},$ (48)
where $C$ is some constant with dimension $[1/l]$ that arises due to the transition from the sum to the integral. We observe that $\frac{\overline{\chi_{\Delta E}}}{2\Delta E}$ is a function that equals $1/2\Delta E$ in the energy region near $E_{K}$ with width $\Delta E$, and is zero elsewhere. As $\Delta E\to 0$, the region of energies shrinks to zero, but the value of the transmission rate grows to infinity. We can see that the integral of $\frac{\overline{\chi_{\Delta E}}}{2\Delta E}$ is a function that equals $1/2\Delta E$ over energy, and its total integral is one. Thus, in the weak limit, the function approaches $\delta(E-E_{k})$. Therefore, for transmissions between neighborhoods of states $k$ and $l$, we obtain:
@@
Here we can see what we have lost by using simplification (27): the limit of $\frac{\chi_{\Delta E}}{2\Delta E}$ at $\Delta E\to 0$ is not just a delta function, but a sum $\delta(E-E_{k})+\text{v.p.}\frac{i}{E-E_{k}}$, where v.p. means that taking an integral over $E$ one should take its principal value. This expression is very well-known in the context of perturbation theory. The second term is responsible for perturbational corrections to the values of energies of states and matrix elements of interaction.
We could have chosen a more rigorous method for deriving equation (30). We would still seek a smoothed unitary evolution equation, which we achieve by expressing the matrix $\mathcal{V}$ in the form of the Dyson series. If we only retain the first term of the series for $\mathcal{V}$, it gives the same expression that we have found. Seemingly, this method allows us to find higher-order terms of the perturbation series, including corrections to energy levels due to interactions and second-order transitions. However, this approach is lengthy, and if we were to take this route, we would aim to investigate perturbation theory within the described framework. In this paper, our primary goal is to deduce the non-coherent evolution equation from the smoothed unitary evolution equation (30). We leave the task of finding the most accurate form of the smoothed unitary evolution equation, which includes perturbative corrections, for subsequent publications.
@@ IV Applications to statistical mechanics
As we mentioned earlier, the equation of non-coherent evolution (40) is a continuous-time Markov chain (CTMC) equation. It is known [42] that such equations are irreversible and converge to a stationary distribution for large values of $t$. The specific distribution depends on the particular $\mathcal{Q}$-matrix. However, the $\mathcal{Q}$-matrix of non-coherent evolution, defined by expressions (36, 38), is symmetric. For symmetric $\mathcal{Q}$-matrices, the stationary distribution is always the same: an equiprobability of all microstates.
@@ V Summary and discussion
Then we move on from studying a discrete series of transformations to continuous evolution. We derive the non-coherent evolution equation, which is the central result of this work (40, 45). Our analysis begins with partitioning the wave function of the system into two components: fast-varying psi functions of microstates and probability amplitudes, which are smooth (in-time) envelopes of the psi function (21). Our derivation relies on the assumption of weak interaction and the existence of well-separated time scales. We introduced a technical timescale $\Delta t$, chosen between the microstate oscillation period and the characteristic scale of amplitude variation. Using this technical assumption, we showed that for timescales much larger than the coherence time, unitary evolution transforms into symmetric stochastic evolution, described by continuous-time Markov chains. We then show that the final results of the calculations do not depend on the value of the timescale $\Delta t$.
@@
We continue by showing that the equation of non-coherent evolution serves as a common basis for both equilibrium statistical mechanics and kinetic theory. We derived the Gibbs microcanonical ensemble as its stationary solution, under the assumption of weak interaction, and showed that this stationary distribution corresponds to the equiprobability of microstates. We also derived the Boltzmann collision integrals by transitioning from equations for probabilities to equations for average values. This derivation justifies the irreversibility of the Boltzmann equation as a direct consequence of the irreversibility of non-coherent evolution. We note that the Boltzmann collision integral is derived strictly from the non-coherent evolution equation and basic definitions, without any ”hand-waving”.
@@
Second, let us turn to quantum thermodynamics. While it is true that the equation of non-coherent evolution (40, 45) is in many respects analogous to those found in quantum thermodynamics, we hope it is clear that both the physical picture offered by the present theory and the mathematical methods employed are very different. Concepts such as coherence length are not introduced in quantum thermodynamics. We may hope that further development of the theory will provide a testable and decisive distinction between the two approaches.
@@
It is only when we impose the condition of non-coherence that the situation changes. The system then reduces to one in which all microstates have definite probabilities but lack definite phases, analogous to photons in the non-coherent experiment (Fig. 2c). The evolution of such a system is governed by the stochastic equation (40, 45), which is irreversible. Thus, in the present viewpoint, it is not interaction but non-coherence that constitutes the true origin of irreversibility.
@@
Transport in nanostructures is a rapidly developing field [46], and the distinction between coherent and non-coherent transmission has already been discussed in relation to interfaces [47, 48]. It is therefore quite possible that some effects consistent with the present theory are already reported in the literature. Otherwise, we may hope to design a realistic experiment to verify the theory in the near future.

### v6

@@ III Non-coherent evolution
We suppose $V_{kl}$ and $P_{l}$ are smooth functions of energy at the point of interest, so in a short interval of values of $l_{N}$ we can approximate them as constants. So summation over $l_{N}$ gives the number of states inside the energy region near llN(l)N(l) (Fig. 4b):
$\frac{\pi}{\hbar\Delta E}\sum_{l_{T},l_{N}}|V_{kl}|^{2}P_{l_{T},l_{N}}=\frac{\pi N(l)}{\hbar\Delta E}\sum_{l_{T}}|V_{k\,l_{T}}|^{2}P_{l_{T}}=\frac{2\pi}{\hbar}\sum_{l_{T}}|V_{k\,l_{T}}|^{2}P_{l_{T}}\nu(l)$ (54)
where $\nu(l)=N(l)/(2\Delta E)$ is the density of states near $l$. Multiplier 2 arises because $\Delta E$ is the absolute value of the deviation of energy from $E_{l}$ and the energy in the region can be less than $E_{l}$ and higher than $E_{l}$ so the length of the region is $2\Delta E$.
We can reason differently. We perform the same subsitution $\Delta t=\pi\hbar/\Delta E$ as in (53), and approximate the sum as an integral (we recall that we have omitted $\overline{\chi}_{kl}$ in the expression for the summation for brevity):
$\frac{1}{\hbar^{2}}\sum_{l}|V_{kl}|^{2}\Delta tP_{l}=\frac{2\pi}{\hbar}C\int dl\frac{\overline{\chi}_{kl}}{\Delta E}|V_{kl}|^{2}P_{l},$ (55)
where $C$ is some constant with dimension $[1/l]$ that arises due to the transition from the sum to the integral. We want to illustrate this transition from summation to integration with an example commonly used in solid-state physics [51]. It is the transition from summation over momenta $p$ to integration over them. We use a one-dimensional example for simplicity: $\sum_{p}\to\frac{L}{2\pi\hbar}\int dp$. Here $L$ is the length of the system under consideration. Typically, it cancels out in the final result. In this example, the constant $C=\frac{L}{2\pi\hbar}$.
We observe that $(\overline{\chi}_{kl})/(2\Delta E)$ is a function that equals $1/(2\Delta E)$ in the energy region near $E_{K}$ with width $\Delta E$, and is zero elsewhere. As $\Delta E\to 0$, the region of energies shrinks to zero, but the value of the transition rate grows to infinity. We can see that the integral of $(\overline{\chi}_{\Delta E})/(2\Delta E)$ is a function that equals $1/(2\Delta E)$ over energy, and its total integral is one. Thus, in the weak limit, the function approaches $\delta(E-E_{k})$. Therefore, for transmissions between neighborhoods of states $k$ and $l$, we obtain:
@@
At this point we can understand what we have gained by changing $\chi_{kl}$ to $\overline{\chi}_{kl}$ (see Eqs. 29, 34). If not for this exchange, we would have to keep the function $\chi_{kl}$ inside the summations over $l$ after Eq. (36). We would have to carry it through all the manipulations concerning non-coherence. The function $\chi_{kl}$ is ”glued” to the interaction $V_{kl}$, so when we get the squares of the latter, it would be combined with the former: $|\chi_{kl}V_{kl}|^{2}$. So in place of expression (57) we would have:
$\frac{1}{\hbar^{2}}\sum_{l}|\chi_{kl}V_{kl}|^{2}\Delta tP_{l}=\frac{2\pi}{\hbar}C\int dl\frac{|\chi_{kl}|^{2}}{\Delta E}|V_{kl}|^{2}P_{l}.$ (57)
Because of condition (31), the subsequent argument holds and we obtain exactly the same result (56). The argument that appeals more to visual intuition and leads to the form of equation (54) does not hold with the much more complicated function $\chi_{kl}$.
While the exchange of $\chi_{kl}$ for $\overline{\chi}_{kl}$ allows us to use shorter formulas in the computation and permits more intuitive reasoning, there is very important information that we lose with this exchange. Before we take the square of $\chi_{kl}$, it behaves qualitatively differently from $\overline{\chi}_{kl}$. The limit of $(\chi_{\Delta E})/(2\Delta E)$ at $\Delta E\to 0$ is not just a delta function, but a sum $\delta(E-E_{k})+\text{v.p.}\frac{i}{E-E_{k}}$, where v.p. means that taking an integral over $E$ one should take its principal value. This expression is very well-known in the context of perturbation theory. The second term is responsible for perturbational corrections to the values of energies of states and matrix elements of interaction. When we interchange $\chi_{kl}$ for $\overline{\chi}_{kl}$, we discard any perturbative effects of interactions.
We could have chosen a more rigorous method for deriving equation (37). We would still seek a smoothed unitary evolution equation, which we achieve by expressing the matrix $\mathcal{V}$ in the form of the Dyson series. If we only retain the first term of the series for $\mathcal{V}$, it gives the same expression that we have found. Seemingly, this method allows us to find higher-order terms of the perturbation series, including corrections to energy levels due to interactions and second-order transitions. However, this approach is lengthy, and if we were to take this route, we would aim to investigate perturbation theory within the described framework.In this paper, our primary goal is to deduce the non-coherent evolution equation from the smoothed unitary evolution equation (37). The interchange of $\chi_{kl}$ for $\overline{\chi}_{kl}$ is a shortcut to achieve this goal. We leave the task of finding the most accurate form of the smoothed unitary evolution equation, which includes perturbative corrections, for subsequent publications.
@@ IV Applications to statistical mechanics
As we mentioned earlier, the equation of non-coherent evolution (48) is a continuous-time Markov chain (CTMC) equation. It is known [48] that such equations are irreversible and converge to a stationary distribution for large values of $t$. The specific distribution depends on the particular $\mathcal{Q}$-matrix. However, the $\mathcal{Q}$-matrix of non-coherent evolution, defined by expressions (44, 46), is symmetric. For symmetric $\mathcal{Q}$-matrices, the stationary distribution is always the same: an equiprobability of all microstates.
@@ V Summary and discussion
Then we move on from studying a discrete series of transformations to continuous evolution. We derive the non-coherent evolution equation, which is the central result of this work (48, 52). Our analysis begins with partitioning the wave function of the system into two components: fast-varying psi functions of microstates and probability amplitudes, which are smooth (in-time) envelopes of the psi function (22). Our derivation relies on the assumption of weak interaction and the existence of well-separated time scales. We introduced a technical timescale $\Delta t$, chosen between the microstate oscillation period and the characteristic scale of amplitude variation. Using this technical assumption, we showed that for timescales much larger than the coherence time, unitary evolution transforms into symmetric stochastic evolution, described by continuous-time Markov chains. We then show that the final results of the calculations do not depend on the value of the timescale $\Delta t$.
@@
We continue by showing that the equation of non-coherent evolution serves as a common basis for both equilibrium statistical mechanics and kinetic theory. We derived the Gibbs microcanonical ensemble as its stationary solution, under the assumption of weak interaction, and showed that this stationary distribution corresponds to the equiprobability of microstates. We also derived the Boltzmann collision integrals in two important cases by transitioning from equations for probabilities to equations for average values. This derivation justifies the irreversibility of the Boltzmann equation as a direct consequence of the irreversibility of non-coherent evolution. We note that the Boltzmann collision integral is derived strictly from the non-coherent evolution equation and basic definitions, without any ”hand-waving”.
@@
Second, let us turn to quantum thermodynamics. While it is true that the equation of non-coherent evolution (48, 52) is in many respects analogous to those found in quantum thermodynamics, we hope it is clear that both the physical picture offered by the present theory and the mathematical methods employed are very different. Concepts such as coherence length are not introduced in quantum thermodynamics. We may hope that further development of the theory will provide a testable and decisive distinction between the two approaches.
@@
It is only when we impose the condition of non-coherence that the situation changes. The system then reduces to one in which all microstates have definite probabilities but lack definite phases, analogous to photons in the non-coherent experiment (Fig. 2c). The evolution of such a system is governed by the stochastic equation (48, 52), which is irreversible. Thus, in the present viewpoint, it is not interaction but non-coherence that constitutes the true origin of irreversibility.
@@
Transport in nanostructures is a rapidly developing field [52, 53, 54], and the distinction between coherent and non-coherent transmission has already been discussed in relation to interfaces [55, 56]. It is therefore quite possible that some effects consistent with the present theory are already reported in the literature. Otherwise, we may hope to design a realistic experiment to verify the theory in the near future.